# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

### v1

@@ 4.3.1 Redshift uncertainty damping
As pointed out in the eBOSS QSO analysis [62, 24, 20], a straightforward approach to model the impact of redshift uncertainty is to introduce an additional Gaussian damping term:
$P_{g,\,s}(k,\mu)=P_{\rm{novir}}(k,\mu)\,F_{\rm{FOG}}(k,\mu)\exp\left[-(k\mu\sigma^{\prime})^{2}\right]$ (4.4)
Here, $P_{\rm{novir}}$ is the non-linear kaiser power spectrum and $F_{\rm{FOG}}$ is finger-of-god (FOG) factor. In the last Gaussian damping term, the parameter $\sigma^{\prime}$ quantifies the scale of redshift uncertainty and is related to the equivalent velocity smearing with $\sigma^{\prime}=\Delta v_{\rm{smear}}*{(1+z)}/{H(z)}$ [62]. Using a Taylor expansion of the Gaussian exponential term:
$\exp\left[-(k\mu\sigma^{\prime})^{2}\right]\,\rightarrow 1+a_{2}(k\mu)^{2}+a_{4}(k\mu)^{4}+\ldots$ (4.5)
We note that the second and third terms scale as $P(k)(k\mu)^{2}$ and $P(k)(k\mu)^{4}$, respectively. These terms show similar structures to those of EFT counterterms in Eq.˜4.3. In the EFT of LSS, the general form of such counterterms can be derived from a moment expansion of the redshift-space power spectrum [51], as well as from comparisons with phenomenological models [6]. These studies suggest that this part of EFT counterterms, parameterized by $\alpha_{2}$ and $\alpha_{4}$ particularly, are capable of absorbing nonlinear effects introduced by redshift uncertainty. We will demonstrate this explicitly in Section˜5 with detailed explanation in Appendix˜B.
@@ 4.3.2 Redshift catastrophics correction
Redshift catastrophics have been addressed in theoretical models for redshift surveys like WiggleZ [14] and recent Euclid [23]. Assuming that the redshift catastrophics contamination rate is $f_{c}$, an approximated correction is to multiply the EFT galaxy power spectrum by a factor of $(1-f_{c})^{2}$:
@@
The validity of this approximation depends on both $f_{c}$ and the typical displacement caused by catastrophics. Based on our mock catalogs study with contamination $f_{c}=1\%$ and $f_{c}=5\%$, we find that the relative difference between the correction factor, $R_{\rm{corr}}=(1-f_{c})^{2}$, and the ratio from the mocks, $R_{\rm{mock}}$, satisfies:
$\Bigg{\lvert}\frac{R_{\rm{corr}}(k)-R_{\rm{mock}}(k)}{R_{\rm{mock}}(k)}\Bigg{\lvert}<10\%,\quad\rm{For}\ k\gtrsim 0.02\,h\,\rm{Mpc}^{-1}$ (4.7)
This suggests that the approximation is sufficiently accurate for full-shape power spectrum analyzes, especially considering that additional nuisance parameters in the EFT can absorb residual mismatches. However, this correction is not applied in the standard EFT full-shape analysis. This is because, as we will demonstrate in Section˜5.2.2, introducing this free parameter $f_{c}$ in the fitting process will lead to the loss of constraining power on certain cosmological parameters.
@@ 4.4 Cosmological Inference and Covariance Matrices
For cosmological inference, Bayesian statistics is employed to infer constraints on physical or cosmological parameters. Assuming the evidence from the data is normalized to one, the posterior probability is expressed as $P(\Theta|\mathcal{D})\propto\mathcal{L}(\mathcal{D}|\Theta)P(\Theta)$, with $\Theta$ being the set of input parameters, including compressed parameters $\left\{\alpha_{\mathrm{iso}},\;\alpha_{\mathrm{ap}},\;df,\;dm\right\}$ by ShapeFit or cosmological parameters $\left\{h,\;\omega_{cdm},\;\omega_{b},\;\rm{ln}(10^{10}A_{s}),\;n_{s}\right\}$ by Full-Modeling. It also contains nuisance parameters considered in the EFT model: $\left\{b_{1},\;b_{2},\;b_{s},\;\alpha_{0},\;\alpha_{2},\;\alpha_{4},\;\rm{SN}_{0},\;\rm{SN}_{2}\right\}$. Assuming that we have Gaussian-distributed data, the likelihood function can be expressed as $\mathcal{L}(\Theta)\propto e^{-\chi^{2}(\Theta)/2}$. The $\chi^{2}$ statistic quantifies the goodness of fitting and is computed based on the difference between the observed data and the model predictions, which reads as:
@@
In our study, the vector $\mathcal{D}$ includes a combination of the power spectrum monopole $P_{0}(k)$ and quadrupole $P_{2}(k)$. The subscript ’data’ and ’model’ refers to the measurements from observations (here mocks) and theoretical model predictions. The covariance matrix $\mathcal{C}$ is extracted from Quijote 500 mock catalogs. It incorporates the variances observed in the data and is normalized to the number of the datasets, $N_{d}$, with the formulation:
$\mathcal{C}_{i,j}=\frac{1}{N_{d}-1}\ \,\langle(\mathcal{D}_{i}-\overline{\mathcal{D}}_{i})(\mathcal{D}_{j}-\overline{\mathcal{D}}_{j})\rangle$ (4.9)
Where $\mathcal{D}_{i}$ signifies the power spectrum from a single realization and $\overline{\mathcal{D}}_{i}$ is ensemble mean. In Quijote suite, the volume of each realization is $V_{1}=1\,h^{-3}\rm{Gpc}^{3}$. To model the reduced cosmic variance expected from galaxy surveys, we use the rescaled covariance by volume as: $\mathcal{C^{\prime}}=({V_{1}}/{V_{n}})\,\mathcal{C}$, where $V_{n}$ is the target survey volume. In our analyses, we rescale the covariance to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$ to match the expected survey volumes of DESI and Euclid.
@@ 5.1 Impact of redshift errors on clustering statistics
Figure 2: Comparison of power spectrum monopole (left) and quadrupole (right) between clean mocks and mocks contaminated by redshift uncertainty. The contaminated mocks are generated to mimic the observational effect of LRG-like and QSO-like smearing. The error bars, $\sigma$, are rescaled to match an effective survey volume of $V_{25}$. The middle subpanels display the power spectrum ratios, $R_{\ell}(k)$, between the contaminated and clean mocks. The lower subpanels show the their residuals in units of $\sigma$, with the dotted line indicating the $2\sigma$ threshold. We see significant clustering deviations ($2\sigma$) due to QSO-like Gaussian or Lorentzian smearing for $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$.
@@
Figure 3: Same plot as Figure˜2 comparing clean and mocks contaminated by ELG-like catastrophics and slitless-like errors. The middle subpanels show that the power spectrum ratios $R(k)$ and expected reduction $(1-f_{c})^{2}$ due to the redshift catastrophics (dash line). Slitless-like errors introduce an overall suppression of power spectrum amplitude by $10\%{-}12\%$, exceeding $3\sigma$ level in both monopole and quarupole.
@@
The impact of redshift errors on the clustering statistics is presented in Figure˜2 and Figure˜3. For redshift uncertainty, it introduces a diffusive effect that systematically suppresses the power spectrum amplitude, with the suppression becoming stronger at smaller scales (i.e., larger $k$). As discussed in Section˜4.3.1, this effect can be modeled by a Gaussian damping term. In the middle subpanels of Figure˜2, we plot the ratios of power spectra, $R(k)$, between the contaminated mocks and clean mocks and we find that $R(k)$ exhibits Gaussian-like profiles, with larger $\Delta v_{\rm{smear}}$ resulting in stronger damping of the power spectrum. For an effective survey volume of $V_{25}$, the LRG-like smearing leads to minimal impacts on the clustering, with power spectrum difference $\Delta P\lesssim 2\sigma$ (see lower subpanels in Figure˜2). For QSO-like redshift uncertainty, the Lorentzian and Gaussian smearing introduce a comparable suppression, with $\Delta P>2\sigma$ for $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$ in both the monopole and quadrupole.
@@
Redshift catastrophics affect only a small fraction of objects but cause significant shifts of their positions along the line-of-sight. In Section˜4.3.2, we show that this distortion leads to an overall reduction on the power spectrum, scaling approximately as $(1-f_{c})^{2}$. As illustrated in Figure˜3, the ELG-like catastrophics with $f_{c}=1\%$ reduce the clustering signal to $\sim 98\%$ of its original amplitude, with deviations below $3\sigma$. For slitless-like errors (i.e., $\sigma_{v}=85.7\,\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$), the power spectrum amplitude from contamianted mocks is reduced to $(1-5\%)^{2}=90.25\%$, i.e., a suppression of $10\%$. An additional deviation from the $(1-f_{c})^{2}$ arises due to the smearing effect introduced by redshift uncertainty. In this case, the overall suppression of slitless-like errors is more significant, exceeding $3\sigma$ for $k\gtrsim 0.02\,h\,\rm{Mpc}^{-1}$ in the monopole and $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$ in the quadrupole. Such deviations, if uncorrected, can bias the cosmological inference. In the following subsections, we describe their influences when running the inference in different cosmological cases: $\Lambda\rm{CDM}$, $w_{0}w_{a}\rm{CDM}$ and $\nu\Lambda\rm{CDM}$.
@@ 5.2 Baseline Fitting
Our baseline full-shape analysis fits the power spectrum monopole and quadrupole over the range $0.02\leq k\leq 0.2\,h\,\rm{Mpc}^{-1}$, using the standard EFT model without applying the correction factor of redshift catastrophics. We adopt two fitting strategies, ShapeFit and Full-Modeling, to extract cosmological information. The results are presented within the standard flat-$\Lambda\rm{CDM}$ model, assuming a BBN prior on baryon abundance $\Omega_{\mathrm{b}}$, and wide Gaussian prior on spectral index $n_{s}$. The cosmological inference is performed using the desilike [^3] https://github.com/cosmodesi/desilike package and parameter priors are set to the same as the those in DESI DR1 full-shape analysis [61]. The parameter space is sampled with the code emcee [63]. All plots and confidence intervals are generated using the GetDist  package [64]. Beside the baseline full-shape analysis, we also evaluate the impact of redshift errors on BAO fitting. These results are discussed in Appendix˜A.
@@ 5.2.1 Cosmological influence
Figure 4: Distributions of the best-fit values for ShapeFit physical parameters. Scatter plots compare results from clean and contaminated mocks in Table˜1, with each color representing a different error type. The corresponding marginalized histograms are shown in the right and top subpanels. Each point represents one fit to the mean of 25 random mocks, with the covariance rescaled to volume $V_{25}$. Overall, the parameter estimates remain consistent across most cases, except for $df$, which shows a notable deviation in the mocks contaminated by slitless-like errors.
@@
Figure 5: Same plots as Figure˜4 for Full-Modeling cosmological parameters. As before, results from clean and contaminated mocks in Table˜1 are compared. Only $\ln(10^{10}A_{s})$ shows a notable deviation in mocks contaminated by slitless-like errors.
@@
In this section, we assess the impact of redshift errors on cosmological inference. These results are based on 200 independent fits to the Quijote mocks catalogs in Table˜1. For each fit, we use the mean power spectrum derived from 25 randomly selected mocks, with a covariance matrix rescaled to the volume $V_{25}$. Comparisons of the best-fit parameters between analyses from clean and contaminated mocks are presented in Figure˜4 for ShapeFit and in Figure˜5 for Full-Modeling. The corresponding $\chi^{2}$ comparison, provided in Appendix˜B, confirms that all our fits achieved comparable goodness-of-fit.
@@
In Figure˜4 and Figure˜5, the upper panels illustrate the impact of redshift uncertainty on cosmological inference. The fitting results of ShapeFit and Full-Modelling indicate that there is no significant shift in parameter estimates (<5%). As discussed in Section˜4.3.1, the suppression of the power spectrum caused by redshift uncertainty is expected to be absorbed by the EFT counterterms. In Appendix˜B, we plot the posterior distributions for all nuissance parameters from a representative fit, which shows negative shifts in counterterm parameters $\alpha_{2}$ and $\alpha_{4}$ for contaminated case. This is consistent with our theory explanations. Additionally, by comparing QSO-like uncertainties modeled with Gaussian and Lorentizan profile, we find that, as long as they have the same magnitude of smearing, the presence of extended tail does not significantly affect cosmological measurements in full-shape analysis.
@@
The lower panels of Figure˜4 and Figure˜5 assess the impact of redshift catastrophics and hypothetical slitless-like errors on cosmological inference. The ELG-like catastrophics with $f_{c}=1\%$ show negligible bias in parameter estimates (with the largest discrepancy $6\%$), consistent with the previous findings in [21]. In contrast, slitless-like errors with $f_{c}=5\%$ show noticeable biases, particularly in estimating the fractional growth rate parameter $df$ in ShapeFit and the log primordial amplitude parameter $\ln(10^{10}A_{s})$ in Full-Modeling. Both parameters are systematically underestimated by $6\%{-}16\%$ compared to clean case. We identify that this negative shift arises from the $5\%$ catastrophics contamination in slitless-like errors, which significantly suppresses the power spectrum amplitude (see Figure˜2). If not adequately modeled in theory, this effect could bias the estimates of amplitude-related parameters. Nevertheless, as discussed in Section˜4.3.2, a simple correction by multiplying the power spectrum by a factor of $(1-f_{c})^{2}$ can mitigate this suppression. In the following section, we incorporate this correction into the cosmological analysis to further investigate the influence of redshift catastrophics.

### v4

@@ 4.3.1 Redshift uncertainty damping
Figure 2: Power spectrum monopole and quadrupole ratios from the mean of 500 mocks (dashed line) compare to theoretical model (solid lines). The shaded regions correspond to the errors from the standard deviation of the 500 mocks divided by 5. Left: Comparison between the ratios from mocks contaminated by Gaussian or Lorentzian smearing (characterized by $\sigma_{\Delta v}$ and $\mathrm{w}_{\Delta v}$ respectively) relative to clean mocks with the Gaussian Damping (GD) or Lorentzian Damping (LD) factor. Right: Comparison between $f_{c}=1\%$ and $5\%$ contamination mock ratios with the $(1-f_{c})^{2}$ scaling.
As pointed out in the eBOSS QSO analyses [smith_completed_2020, hou_completed_2020], an approach to model the impact of redshift uncertainty in power spectrum is to introduce an additional damping term $D(k,\mu)$:
$P_{g,\,s}^{\rm{smear}}(k,\mu)=P_{g,\,s}(k,\mu)\,D(k,\mu)$ (4.4)
$D(k,\mu)=\begin{cases}\exp\!\left[-(k\mu\sigma_{\rm{zerr}})^{2}\right],&\text{Gaussian damping},\\ 1/\left[1+(k\mu\mathrm{w}_{\rm{zerr}})^{2}\right],&\text{Lorentzian damping}.\end{cases}$ (4.5)
where the redshift error parameters $\sigma_{\rm{zerr}}$ or $\mathrm{w}_{\rm{zerr}}$ quantify the scale of redshift uncertainty. For the Gaussian smearing case, $\sigma_{\rm{zerr}}$ is related to the smearing value with $\sigma_{\rm{zerr}}=\sigma_{\Delta v}*{(1+z)}/{H(z)}$, and, as shown in the left panels of Figure˜2, this smearing effects are well captured by the Gaussian damping factor up to $k\sim 0.3\,h\,\rm{Mpc}^{-1}$. In the Lorentzian smearing case, the parametrization implies an additional factor of two, i.e. $\mathrm{w}_{\rm{zerr}}=2\,\mathrm{w}_{\Delta v}*{(1+z)}/{H(z)}$. With Lorentzian damping, we see deviations compared to our mock analyses (orange lines in Figure˜2). This is possibly due to the velocity cut we applied when constructing the Lorentzian smearing profile. However, as we will demonstrate in the following, these deviations can be effectively absorbed within the more flexible EFT framework.
To compare the damping factor with EFT model, we can apply the Taylor expansion on the damping terms. For both Gaussian and Lorentzian smearing, we have:
$D(k,\mu)\,\rightarrow 1+a_{2}(k\mu)^{2}+a_{4}(k\mu)^{4}+\ldots$ (4.6)
We notice that the second and third terms scale as $P(k)(k\mu)^{2}$ and $P(k)(k\mu)^{4}$, respectively. These terms have similar structures to those of EFT counterterms in Eq.˜4.3. In the EFT of LSS, the general form of such counterterms can be derived from a moment expansion of the redshift-space power spectrum [senatore_redshift_2014, aviles_redshift_2021, damico_taming_2024], as well as from comparisons with phenomenological models [ivanov_cosmological_2020]. Consistently, [simon_cosmological_2023] has shown that the leading corrections coming from redshift uncertainty are degenerate with EFT counterterms, particularly those parameterized by $\alpha_{2}$ and $\alpha_{4}$. In section˜5.2.1, we demonstrate equivalently that the redshift uncertainty effect can be absorbed by the EFT nuisance parameters. Therefore, when we do the cosmological fitting, we use the EFT-based clustering model without including the additional damping factor.
@@ 4.3.2 Redshift catastrophics correction
Redshift catastrophics have been addressed in theoretical models for redshift surveys like WiggleZ [blake_wigglez_2010] and more recently Euclid [euclid_collaboration_euclid_2025-5]. Assuming that the redshift catastrophic contamination rate is $f_{c}$, an approximate correction is to multiply the EFT galaxy power spectrum by a factor of $(1-f_{c})^{2}$:
@@
The validity of this approximation depends both on $f_{c}$ and the displacement caused by catastrophics. In Figure˜2, based on our mock analyses with contamination $f_{c}=1\%$ and $5\%$, we find that the relative difference between the correction factor, $R_{\rm{corr}}\equiv(1-f_{c})^{2}$, and the ratio from the mock statistics, $R_{\rm{mock}}\equiv P_{\text{dv},\,f_{c}}/P_{\rm{clean}}$, satisfies:
$\Bigg\lvert\frac{R_{\rm{corr}}(k)-R_{\rm{mock}}(k)}{R_{\rm{mock}}(k)}\Bigg\rvert\lesssim 6\%,\quad\rm{For}\ 0.02\leq k\leq 0.3\,h\,\rm{Mpc}^{-1}$ (4.8)
This suggests that the approximation is sufficiently accurate for full-shape power spectrum analysis, especially considering that additional nuisance parameters in the EFT can absorb residual mismatches. In DESI, the $f_{c}$ is typically below 1%. Their overall impact is expected to be small [yu_elg_2025] and such correction is not included in the standard EFT model. However, in slitless surveys such as Euclid, where $f_{c}$ is significantly higher, we will show in section˜5.2.2 that incorporating this correction will be necessary to obtain unbiased cosmological constraints.
@@ 4.4 Cosmological inference and Covariance matrices
Cosmological inference employs Bayesian statistics to infer constraints on physical or cosmological parameters. Assuming the evidence from the data is normalized to one, the posterior probability is expressed as $P(\Theta|\mathcal{D})\propto\mathcal{L}(\mathcal{D}|\Theta)P(\Theta)$, with $\Theta$ being the set of input parameters, including compressed parameters $\left\{\alpha_{\mathrm{iso}},\;\alpha_{\mathrm{ap}},\;df,\;dm\right\}$ by ShapeFit or cosmological parameters $\left\{h,\;\omega_{cdm},\;\omega_{b},\;\rm{ln}(10^{10}A_{s}),\;n_{s}\right\}$ by Full-Modeling. It also contains nuisance parameters considered in the EFT model: $\left\{b_{1},\;b_{2},\;b_{s},\;\alpha_{0},\;\alpha_{2},\;\alpha_{4},\;\rm{SN}_{0},\;\rm{SN}_{2}\right\}$. Assuming that we have Gaussian-distributed data, the likelihood function can be expressed as $\mathcal{L}(\Theta)\propto e^{-\chi^{2}(\Theta)/2}$. The $\chi^{2}$ statistic quantifies the goodness of fitting and is computed based on the difference between the observed data and the model predictions, which reads as:
@@
In our study, the vector $\mathcal{D}$ includes a combination of the power spectrum monopole $P_{0}(k)$ and quadrupole $P_{2}(k)$. The subscript ’data’ and ’model’ refers to the measurements from observations (here mocks) and theoretical model predictions. For each type of redshift-error, we have constructed 500 corresponding contaminated mocks and use them to build the data vector and covariance matrix. The covariance matrix $\mathcal{C}$ is computed and normalized to the number of the datasets, $N_{d}$, with the formulation:
$\mathcal{C}_{ij}=\frac{1}{N_{d}-1}\sum_{n=1}^{N_{d}}\left(\mathcal{D}_{i}^{(n)}-\overline{\mathcal{D}}_{i}\right)\left(\mathcal{D}_{j}^{(n)}-\overline{\mathcal{D}}_{j}\right),$ (4.10)
The $\mathcal{D}_{i}^{(n)}$ denotes the $i$-th bin of the data vector from the $n$-th mock realization and $\overline{\mathcal{D}}_{i}$ is the ensemble mean. In Quijote suite, the volume of each realization is $V_{1}=1\,h^{-3}\rm{Gpc}^{3}$. To model the reduced cosmic variance expected from galaxy surveys, we use the rescaled covariance as: $\mathcal{C^{\prime}}=({V_{1}}/{V_{n}})\,\mathcal{C}$, where $V_{n}$ is the target survey volume. In our analyses, we rescale the covariance to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$ to match the expected survey volumes of DESI and Euclid.
@@ 5.1 Impact of redshift errors on clustering statistics
Figure 3: Comparison of power spectrum monopole (left) and quadrupole (right) between clean mocks and mocks contaminated by redshift uncertainties. The error bars, $\sigma$, are rescaled to match an effective survey volume of $V_{25}$. The lower panels show their residuals in units of clean mocks $\sigma$, with the dotted line indicating the $3\sigma$ threshold. In QSO-like Gaussian or Lorentzian smearing cases, we see significant deviations ($3\sigma$) for $k\gtrsim 0.09\,h\,\rm{Mpc}^{-1}$.
@@
Figure 4: Similar to Figure˜3 comparing clean mocks and mocks contaminated by ELG-like catastrophics and slitless-like errors. Redshift catastrophics introduce an overall suppression on power spectrum amplitude, leading $3\sigma$ difference for sliltess-like errors.
@@
The impact of redshift errors on the clustering statistics is presented in Figure˜3 and Figure˜4. For redshift uncertainty, it introduces a smearing effect that systematically suppresses the power spectrum amplitude, with the suppression becoming stronger at smaller scales (i.e., larger $k$). The lower panels of Figure˜3 show the residual plot in units of the clean mocks error bar $\sigma$. For an effective survey volume of $V_{25}$, the LRG-like smearing leads to minimal impacts on the clustering, with power spectrum difference $\Delta P<3\sigma$. For QSO-like redshift uncertainty, both Lorentzian and Gaussian smearing lead to $\Delta P>3\sigma$ for $k\gtrsim 0.09\,h\,\rm{Mpc}^{-1}$ in the monopole and quadrupole. Notably, the QSO Lorentzian smearing, with the same RMS value as the Gaussian case, shows a comparatively flatter suppression at large $k$.
@@
Redshift catastrophics affect only a small fraction of objects but cause significant shifts on their positions along the line-of-sight. Section˜4.3.2 shows that this distortion leads to an overall reduction on the power spectrum, scaling approximately as $(1-f_{c})^{2}$. As illustrated in Figure˜4, the ELG-like catastrophics with $f_{c}=1\%$ leads to a 2% reduction of the clustering amplitude, with maximum deviations around $3\sigma$. However, slitless-like errors (i.e., $\sigma_{v}=85.7\,\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$) introduce an overall suppression of the power spectrum amplitude by 10%-12%, resulting in a deviation $3\sigma$ for $k\gtrsim 0.02\,h\,\rm{Mpc}^{-1}$ in the monopole and $k\gtrsim 0.12\,h\,\rm{Mpc}^{-1}$ in the quadrupole. Such $3\sigma$ deviations, if uncorrected, could bias cosmological inferences. In the following subsections, we will describe their impact when running the inference in different cosmological cases: $\Lambda\rm{CDM}$, $w_{0}w_{a}\rm{CDM}$ and massive neutrino.
@@ 5.2 Baseline Fitting
Our baseline full-shape analysis fits the power spectrum monopole and quadrupole over the range $0.02\leq k\leq 0.2\,h\,\rm{Mpc}^{-1}$, using the standard EFT model without including the damping term from redshift uncertainty or the correction factor from redshift catastrophics. We adopt two fitting strategies, ShapeFit and Full-Modeling, to extract cosmological information. The results are presented within the standard flat-$\Lambda\rm{CDM}$ model assuming a BBN prior on baryon abundance $\Omega_{\mathrm{b}}$ and wide Gaussian prior on spectral index $n_{s}$. The cosmological inference is performed using the desilike [^3] https://github.com/cosmodesi/desilike package and parameter priors are set to the same as the those in DESI DR1 full-shape analysis [desi_collaboration_desi_2025-1]. The parameter space is sampled with the code emcee [foreman-mackey_emcee_2013]. All plots and confidence intervals are generated using the GetDist  package [lewis_getdist_2025]. Beside the baseline full-shape analysis, we also evaluate the impact of redshift errors on BAO fitting. These results are discussed in Appendix˜A.
@@ 5.2.1 Cosmological influence
Figure 5: Distributions of the best-fit values for ShapeFit compressed parameters. Scatter plots compare results from clean and contaminated mocks in Table˜1. Each point represents one fit to the mean of 25 random mocks, with the covariance rescaled to volume $V_{25}$. The corresponding marginalized histograms are shown in the right and top subpanels. The parameter estimates remain consistent across most cases, except for dfdf, which shows a notable deviation in mocks contaminated by slitless-like errors.
@@
Figure 6: Same plots as Figure˜5 for Full-Modeling cosmological parameters. As before, results from clean and contaminated mocks in Table˜1 are compared. Only $\ln(10^{10}A_{s})$ shows a notable deviation in mocks contaminated by slitless-like errors.
Figure 7: Posterior distribution for nuisance parameters in our baseline EFT fitting. Constraints are compared from clean mocks (dark blue) and those contaminated by redshift uncertainties: LRG-like Gaussian (purple), QSO-like Gaussian (pink) and QSO-like Lorentzian (red) smearings. Contours show the 68% and 95% confidence level. Results are derived from fits to the mean power spectrum monopole and quadrupole of 25 mocks using covariance rescaled to $V_{25}$. The marginalized posteriors of $\alpha_{2}$ and $\alpha_{4}$ show negative shifts, reflecting the smearing effect of redshift uncertainty, while the positive shift of $\mathrm{SN}_{2}$ shows the difference brought by the Lorentzian smearing.
@@
In this section, we assess the impact of redshift errors on cosmological inference. These results are based on 200 independent fits to the Quijote mocks catalogs in Table˜1. For each fit, we use the mean power spectrum derived from 25 randomly selected mocks (drawn from 500), with a covariance matrix rescaled to the volume $V_{25}$. Comparisons of the best-fit parameters between analyses from clean and contaminated mocks are presented in Figure˜5 for ShapeFit and in Figure˜6 for Full-Modeling. The corresponding $\chi^{2}$ comparison, provided in Appendix˜B, confirms that all our fits achieved comparable goodness-of-fit.
@@
In Figure˜5 and Figure˜6, the upper panels illustrate the impact of redshift uncertainty on cosmological inference. The fitting results from both ShapeFit and Full-Modelling indicate that no significant bias (<5%) in cosmological parameter estimates. As discussed in section˜4.3.1, the suppression of the power spectrum caused by redshift uncertainty is expected to be absorbed by EFT counterterms. Figure˜7 plots the posterior distributions of all nuisance parameters. These posteriors reveal negative shifts in $\alpha_{2}$ and $\alpha_{4}$, consistent with the expected absorption of the damping effect. Another parameter that shows variation is $\mathrm{SN}_{2}$, which represents the second-order stochastic term and scales as $(k\mu)^{2}$. Its positive shift in the Lorentzian smearing case could be led by the flatter suppression pattern of the Lorentzian profile at large $k$.
@@
The lower panels of Figure˜5 and Figure˜6 assess the impact of redshift catastrophics and hypothetical slitless-like errors on cosmological inference. The ELG-like catastrophics with $f_{c}=1\%$ show negligible bias in parameter estimates (with the largest discrepancy $6\%$), consistent with the previous findings in [yu_elg_2025]. In contrast, slitless-like errors with $f_{c}=5\%$ show noticeable biases, particularly in estimating the fractional growth rate parameter $df$ in ShapeFit and the log primordial amplitude parameter $\ln(10^{10}A_{s})$ in Full-Modeling. Both parameters are systematically underestimated by $6\%{-}16\%$ compared to clean case. We identify that this negative shift arises mainly from the $5\%$ catastrophic contamination in slitless-like errors, which significantly suppresses the power spectrum amplitude (see Figure˜3). If not adequately modeled in theory, this effect could bias the estimates of amplitude-related parameters. Nevertheless, in section˜4.3.2, an approximate correction by multiplying the power spectrum by a factor of $(1-f_{c})^{2}$ can mitigate this suppression. In the following section, we incorporate this correction into the cosmological analyses to further investigate the influence of redshift catastrophics.